# Large $N$ Wess-Zumino model at finite temperature and large chemical potential in $3d$

### v2

@@ 1 Introduction
$\lim_{\mu\rightarrow\infty}a^{b}_{\cal O}[0,l]=\frac{\mu^{l+1}\Gamma(l+1)}{2^{2-l}\pi\Gamma(2l+1)}\Bigg{[}1+\frac{\sqrt{4\lambda^{2}+3\pi^{2}}-2\lambda}{6\lambda}+$ $\sum_{n=1}^{l}\frac{(l-n+1)_{2n}\zeta(n+1)}{2^{n}n!}\frac{T^{n+1}}{|\mu|^{n+1}}\Bigg{]}.$ (4)
The corrections appear as $O(e^{-\frac{\mu}{T}})$. The above expression also contains overall constant factors along with $\langle{\cal O}[0,l]\rangle$ due to the OPE coefficient which does not depend on the dynamics at finite temperature. And at large $\mu$ the fermion bilinear operator ${\cal O}_{+}[0,l]$ has thermal expectation given by the following expansion
@@ 2.1 Scalar two point function
${\rm Traces}=\frac{2}{(l+1)(d+l-1)}\left(\delta^{\mu\mu_{1}}\gamma^{\rho\;\alpha\beta}{\cal O}^{\rho_{\mu_{2}}\cdots\mu_{l}}_{\beta\alpha}+\delta^{\mu_{1}\mu_{2}}\gamma^{\rho\;\alpha\beta}{\cal O}^{\rho_{\mu_{3}}\cdots\mu_{l}\mu_{1}}_{\beta\alpha}+{\rm cyclic}\right).$ (14)
@@
Thus combining all the three terms in the OPE (2.1) as obtained in (11), (2.1) and (2.1) respectively, and if the anomalous dimensions of the operators present in the OPE are vanishing at the leading order in large $N$ [^1] Our analysis in the section 4 will establish the fact that for large $N$ Wess-Zumino model in $3d$ operator dimensions do not get corrected at the leading order in large $N$., we can use
@@ 3 Wess-Zumino model with chemical potential($\mu$) and large $\mu$
[…]
@@
where
$\log Z_{b}(m_{b},-i\hat{\mu})$ $=-\sum_{n=-\infty}^{\infty}\int\frac{d^{2}p}{(2\pi)^{2}}\log[\big{(}\frac{2\pi n}{\beta}-\hat{\mu}\big{)}^{2}+p^{2}+m_{b}^{2}],$
$=\frac{1}{6\pi\beta^{2}}\Big{(}\beta^{3}m_{b}^{3}+3\beta m_{b}(\text{Li}_{2}(e^{-\beta(m_{b}+i\hat{\mu})})+\text{Li}_{2}(e^{-\beta(m_{b}-i\hat{\mu})}))$
$\qquad+3\text{Li}_{3}(e^{-\beta(m_{b}+i\hat{\mu})})+3\text{Li}_{3}(e^{-\beta(m_{b}-i\hat{\mu})})\Big{)},$ (46)
@@
$\log Z_{f}(m_{f},-i\hat{\mu})$ $=\sum_{n=-\infty}^{\infty}\int\frac{d^{2}p}{(2\pi)^{2}}\log[\big{(}\frac{(2n+1)\pi}{\beta}-\hat{\mu}\big{)}^{2}+p^{2}+m_{f}^{2}],$
$=-\frac{1}{6\pi\beta^{2}}\Big{(}\beta^{3}m_{f}^{3}+3\beta m_{f}(\text{Li}_{2}(-e^{-\beta(m_{f}+i\hat{\mu})})+\text{Li}_{2}(-e^{-\beta(m_{f}-i\hat{\mu})}))$
$\qquad+3\text{Li}_{3}(-e^{-\beta(m_{f}+i\hat{\mu})})+3\text{Li}_{3}(-e^{-\beta(m_{f}-i\hat{\mu})})\Big{)}.$ (47)
@@
The above pair of coupled transcendental equations are studied numerically at $\mu=0$ in DeWolfe:2019etx . But, at large $\mu$ these equations become analytically tractable and solved in the following manner. Let us first investigate the asymptotic form of the 1st equation from (3) at large $\mu$ as given below, note that as we are looking for real and positive solutions of both $m_{b}$ and $m_{f}$, and we need $10<m_{b}-\mu<$ 1 to have consistent solution for $m_{b}$,
@@ 4.1 Scalar two point function
[…]
@@
×$\times\cos\left(m_{\rm th}\sqrt{(z\mp m)(\pm m-\bar{z})}\right).$ (66)
@@
[…]
@@
[…]
@@
$=\frac{K_{l}}{l!}\sum_{n=0}^{l}\frac{m_{b}^{l-n}(l-n+1)_{2n}}{2^{l+n}n!}\times\Big{(}\text{Li}_{n+1}(e^{-m_{b}-i\hat{\mu}})+(-1)^{l}\text{Li}_{n+1}(e^{-m_{b}+i\hat{\mu}})\Big{)},$ (73)
@@ 4.2 Fermion two point functions
[…]
@@
[…]
@@
$a^{f}_{\mathcal{\tilde{O}}_{+}}[n=0,l]$ $=\frac{l}{{\pi\big{(}\frac{1}{2}\big{)}_{l}}}{\,\sum_{n=0}^{l}\frac{m_{f}^{l-n}(l-n+1)_{2n}}{2^{n+l+2}n!}}\left(\text{Li}_{n+1}(-e^{-i\hat{\mu}-m_{f}})+(-1)^{l}\text{Li}_{n+1}(-e^{-m_{f}+i\hat{\mu}})\right),$ (87)
@@
Applying the inversion formula on the correlator $g_{2}(z,\bar{z})$ we also have obtained the following result in the equation (4.33) of David:2023uya with $d=3$(or $k=1$ as the results obtained there in odd space time dimensions $d=2k+1,\ k\in\mathbb{Z}$),
@@
$=\frac{(l+2)}{2\pi(3+2l)(\frac{1}{2})_{l}}\left(\frac{m_{f}}{2}\right)^{l+2}\sum_{n=0}^{l}\frac{(l+1-n)_{2n}}{(2m_{f})^{n}n!}({\rm Li}_{n+1}(-e^{-m_{f}-i\hat{\mu}})+(-1)^{l}{\rm Li}_{n+1}(-e^{-m_{f}+i\hat{\mu}})),$
@@
Again by using $C_{\psi\psi\cal O}=0$ from (83) the entire analysis of the inversion formula on $g_{3}(x)$ reduces to the computation of the thermal one point functions of fermion bilinears from the $g_{3}(x)$ for the large $N$ Gross-Neveu model performed in David:2023uya , except the fact that one has to keep track of the $e^{im\mu}$ in the correlator (90) compared to the the correlator given in equation (4.33) of David:2023uya with $d=3$(or $k=1$ as the notation used there $d=2k+1$) [^4] Note that there was a typo in the equation (4.33) of David:2023uya , in the denominator of this formula $(k+\frac{1}{2})_{l}$ is not correct, it should be replaced by $(k-\frac{1}{2})_{l}$.. And keeping this $e^{im\mu}$ in that calculation we get
@@
$=\frac{1}{\pi(\frac{1}{2})_{l}}\Big{(}\frac{m_{f}}{2}\Big{)}^{l+2}\sum_{n=0}^{l}\frac{(l-n+1)_{2n}}{(2m_{f})^{n}n!}({\rm Li}_{n+1}(-e^{-m_{f}-i\hat{\mu}})+(-1)^{l}{\rm Li}_{n+1}(-e^{-m_{f}+i\hat{\mu}})).$ (91)
@@
For the one point function of the fermion bilinear ${\cal O}_{0}[0,0]$, the arc contribution has also been evaluated in David:2023uya and it is given by  [^6] See equation (4.49) in David:2023uya , note that the presence of the chemical potential $-i\hat{\mu}$ doesn’t alter this answer, we kept track of an overall -ve sign in our correlator $g_{1}(z,\bar{z})$ in (79) compared to that used in David:2023uya .
@@ 6 Agreement with the stress tensor and spin-1 current
$\log Z$ $=-N\beta V(\frac{m_{f}^{3}}{8\lambda}-\frac{m_{b}^{2}m_{f}}{8\lambda}-\frac{1}{\beta}\log Z_{b}(m_{b},-i\hat{\mu})-\frac{1}{\beta}\log Z_{f}(m_{f},-i\hat{\mu})).$ (117)
@@
Thus the combination $-4a_{\mathcal{O}}[0,2]+2a_{\mathcal{O}_{+}}[0,2]$ agrees with the stress tensor for our model at the free CFT point where the coupling constant $\lambda=0$. We can also see that the expression (6) exactly matches with the stress tensor of the theory (6) at the interacting CFT obtained from the direct calculation of the partition function at $\beta=1$.
@@
$=\frac{1}{{2\pi\beta^{2}}}\Big{[}\beta m_{b}(\log(1-e^{-\beta m_{b}+i\beta\hat{\mu}})-\log(1-e^{-\beta(m_{b}+i\hat{\mu})}))$
$+\text{Li}_{2}(e^{-\beta(m_{b}+i\hat{\mu})})-\text{Li}_{2}(e^{i\beta\hat{\mu}-m_{b}\beta})+\beta m_{f}(\log(1+e^{-\beta(m_{f}+i\hat{\mu})})-\log(1+e^{-\beta m_{f}+i\beta\hat{\mu}}))$
@@ 7 Conclusions
We have studied the large $N$ Wess-Zumino model in presence of the chemical potential($\mu$) in $3d$. Considering the Wess-Zumino model for $2N$ real scalars and $2N$ real fermions without chemical potential in the fundamental of $O(2N)$, we construct the $U(N)$ Wess-Zumino model with $N$ complex scalars and $N$ Dirac fermions. We have introduced a chemical potential in this setup at finite temperature. The non-trivial fixed point of the model is obtained at the saddle point of the action in the zero modes of the auxiliary fields used to linearise the action. The saddle point condition is known as the gap equation. We have found the analytically solvable regime of the gap equations in large $N$ Wess-Zumino model in $3d$ at large chemical potential($\mu$) for arbitrary coupling($\lambda$). By using these analytic solutions of the gap equation at large $\mu$, we evaluate the partition function, stress tensor and spin-1 current of the theory as a perturbative expansion at large $\mu$. Applying the OPE inversion formula we also evaluate the higher spin conserved currents of the theory as a perturbative expansion at large $\mu$. As a direct consequence of the study of the OPE inversion formula we find that at the leading order in large $N$ the OPE coefficients $C_{\phi\phi{\cal O}_{0}[0,l]}$ and $C_{\psi\psi{\cal O}[0,l]}$ vanish for all values of $l$ and thus the relevant three point functions also vanish. Although our OPE analysis can not directly conclude whether the coefficients $C_{\phi\phi{\cal O}_{+}[0,l]}$ and $C_{\phi\phi{\cal O}_{-}[0,l]}$ also vanish. But comparing the results obtained from the partition function we have shown that the $C_{\phi\phi{\cal O}_{+}[0,l]}$ indeed vanishes for $l=1,2$. It would be interesting to evaluate these coefficients directly. An interesting problem would be to extend this study on curved backgrounds such as $S^{1}\times S^{2}$, method developed in David:2024pir for $O(N)$ model in this geometry might be useful. The large chemical potential limit may serve as solvable limits in various other related models.

### v4

@@ 1 Introduction
$\lim_{\mu\rightarrow\infty}a_{\cal O}[0,l]=\frac{\mu^{l+1}\Gamma(l+1)}{2^{2-l}\pi\Gamma(2l+1)}\Bigg{[}1+\frac{\sqrt{4\lambda^{2}+3\pi^{2}}-2\lambda}{6\lambda}+$ $\sum_{n=1}^{l}\frac{(l-n+1)_{2n}\zeta(n+1)}{2^{n}n!}\frac{T^{n+1}}{|\mu|^{n+1}}\Bigg{]}.$ (4)
The corrections appear as $O(e^{-\frac{\mu}{T}})$. The above expression also contains overall constant factors along with $\langle{\cal O}[0,l]\rangle$ due to the OPE coefficient which does not depend on the dynamics at finite temperature. And at large $\mu$ the fermion bilinear operator ${\cal O}_{+}[0,l]$ has thermal expectation given by the following expansion in terms of regularised hypergeometric functions
@@ 2.1 Scalar two point function
${\rm Traces}=\frac{2}{(l+1)(d+l-1)}\left(\delta^{\nu\mu_{1}}\gamma^{\rho\;\alpha\beta}{\cal O}^{\rho{\mu_{2}}\cdots\mu_{l}}_{\beta\alpha}+\delta^{\nu\mu_{2}}\gamma^{\rho\;\alpha\beta}{\cal O}^{\rho{\mu_{3}}\cdots\mu_{l}\mu_{1}}_{\beta\alpha}+{\rm cyclic}\right).$ (14)
@@
Thus combining all the three terms in the OPE (2.1) as obtained in (11), (2.1) and (2.1) respectively, and if the anomalous dimensions of the operators present in the OPE are vanishing at the leading order in large NN [^1] Our analysis in the section 4 will establish the fact that for large $N$ Wess-Zumino model in $3d$ operator dimensions do not get corrected at the leading order in large $N$., we can use
@@ 3 Wess-Zumino model with chemical potential($\mu$) and large $\mu$
The critical Wess-Zumino model is a well-studied example of a CFT at arbitrary coupling strength. The model admits simplifications at large $N$ limit. The non-trivial fixed point of the theory at finite temperature is characterised by the bosonic and fermionic thermal masses denoted by $m_{b}$ and $m_{f}$ respectively, which satisfy the large $N$ gap equations or the saddle point equation of the partition function on $S^{1}\times R^{2}$. We consider the supersymmetric action Moshe:2003xn ; DeWolfe:2019etx for the Wess-Zumino model in $(2+1)d$ in the fundamental of $O(2N)$, with $2N$ real scalars and $2N$ real fermions. Now by combining pairs of real scalars and pairs of real fermions as complex scalars and Dirac fermions respectively we obtain the $U(N)$ Wess-Zumino model with $N$ complex scalars and $N$ Dirac fermions in d=3d=3.222We thank Justin David for suggesting this point.. Now we introduce a chemical potential $\mu$ in this setup. The Euclidean action for $U(N)$ Wess-Zumino model with $N$ complex scalars $\tilde{\phi}_{a}$ and fermions $\tilde{\psi}_{a}$ with sextic intra-species interaction among the complex scalars and the scalar-fermion interation is described below,
@@
$\log Z$ $=\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}d\zeta_{0}d\sigma_{0}e^{-N\beta V(\frac{\sigma_{0}^{3}}{8\lambda}-\frac{\zeta_{0}\sigma_{0}}{8\lambda}-\frac{1}{\beta}\log Z_{b}(\sqrt{\zeta_{0}},\mu)-\frac{1}{\beta}\log Z_{f}(\sigma_{0},\mu))},$ (45)
where
@@
$)\displaystyle=\frac{1}{6\pi\beta^{2}}\Big{(}\beta^{3}m_{b}^{3}+3\beta m_{b}(\text{Li}_{2}(e^{-\beta(m_{b}+i\hat{\mu})})+\text{Li}_{2}(e^{-\beta(m_{b}-i\hat{\mu})}$))
$),\displaystyle\qquad+3\text{Li}_{3}(e^{-\beta(m_{b}+i\hat{\mu})})+3\text{Li}_{3}(e^{-\beta(m_{b}-i\hat{\mu})}$$\Big{)}$, (46)
@@
[…]
$)\displaystyle=-\frac{1}{6\pi\beta^{2}}\Big{(}\beta^{3}m_{f}^{3}+3\beta m_{f}(\text{Li}_{2}(-e^{-\beta(m_{f}+i\hat{\mu})})+\text{Li}_{2}(-e^{-\beta(m_{f}-i\hat{\mu})}$))
$).\displaystyle\qquad+3\text{Li}_{3}(-e^{-\beta(m_{f}+i\hat{\mu})})+3\text{Li}_{3}(-e^{-\beta(m_{f}-i\hat{\mu})}$$\Big{)}$. (47)
@@
The above pair of coupled transcendental equations are studied numerically at $\mu=0$ in DeWolfe:2019etx . But, at large $\mu$ these equations become analytically tractable and solved in the following manner. Let us first investigate the asymptotic form of the 1st equation from (3) at large $\mu$ as given below, note that as we are looking for real and positive solutions of both $m_{b}$ and $m_{f}$, and we need $0<m_{b}-\mu<1$ to have consistent solution for $m_{b}$,
@@ 4.1 Scalar two point function
$\text{where,}\quad K_{l}=\frac{\Gamma(l+1)\Gamma(\nu)}{4\pi\Gamma(l+\nu)},\qquad F_{l}(w)=w^{l+2\nu}{}_{2}F_{1}(l+2\nu,\nu,l+\nu+1,w^{2}).$
@@
$\times\cos\left(m_{\rm th}\sqrt{(z\mp m)(\pm m-\bar{z})}\right).$ (66)
@@
$\hat{a}^{+}_{\rm disc}(\Delta,l)=-\frac{K_{l}}{2\pi}\sum_{m=1}^{\infty}\int_{0}^{1}d\bar{z}\int_{1}^{{\rm max}(1,\frac{1}{m\sqrt{\bar{z}}})}dz\frac{(\bar{z}-1)e^{-i\mu m}F_{l}(\sqrt{\bar{z}})\cos\left(mm_{\rm th}\sqrt{(1-z)(z\bar{z}-1)}\right)}{\bar{z}\left(m^{2}z^{2}\bar{z}\right)^{\frac{\Delta}{2}}\sqrt{(1-z)(z\bar{z}-1)}}.$ (69)
@@
$\hat{a}_{\rm disc}^{+(0)}(\Delta,l)$ $=\frac{2K_{l}}{\pi}\sum_{m=1}^{\infty}\frac{e^{-im\hat{\mu}}}{m^{\Delta}(-\Delta+l+1)}\int_{0}^{\infty}dy\frac{\cos(mm_{\rm th}y)}{(1+y^{2})^{\Delta}},$
@@
$),\displaystyle=\frac{K_{l}}{l!}\sum_{n=0}^{l}\frac{m_{b}^{l-n}(l-n+1)_{2n}}{2^{l+n}n!}\times\Big{(}\text{Li}_{n+1}(e^{-m_{b}-i\hat{\mu}})+(-1)^{l}\text{Li}_{n+1}(e^{-m_{b}+i\hat{\mu}}$$\Big{)}$, (73)
@@ 4.2 Fermion two point functions
$g_{1}(x)=-m_{f}\sum_{\omega_{m}\in{(2m+1)\pi}}e^{-i\hat{\mu}\tau}\int d\omega\delta(\omega-\omega_{m})\int\frac{d^{2}k}{(2\pi)^{2}}\frac{e^{i\omega\tau}e^{i\vec{k}\cdot\vec{x}}}{(\omega-\hat{\mu})^{2}+\vec{k}^{2}+m_{f}^{2}},$ (76)
@@
$g_{1}(x)=-m_{f}\sum_{m\in\mathbb{Z}}\frac{(-1)^{m}e^{im\hat{\mu}}}{(2\pi)^{3}}\int d\omega d^{2}k\frac{e^{i\omega(\tau+m)}e^{i\vec{k}\cdot\vec{x}}}{\omega^{2}+\vec{k}^{2}+m_{f}^{2}}.$ (78)
@@
[…]
@@
Applying the inversion formula on the correlator $g_{2}(z,\bar{z})$ we also have obtained the following result in the equation (4.33) of David:2023uya with d=3d=3(or $k=1$ as the results obtained there in odd space time dimensions $d=2k+1,\ k\in\mathbb{Z}$),
@@
$),\displaystyle=\frac{(l+2)}{2\pi(3+2l)(\frac{1}{2})_{l}}\left(\frac{m_{f}}{2}\right)^{l+2}\sum_{n=0}^{l}\frac{(l+1-n)_{2n}}{(2m_{f})^{n}n!}({\rm Li}_{n+1}(-e^{-m_{f}-i\hat{\mu}})+(-1)^{l}{\rm Li}_{n+1}(-e^{-m_{f}+i\hat{\mu}}$)),
@@
Again by using $C_{\psi\psi\cal O}=0$ from (83) the entire analysis of the inversion formula on $g_{3}(x)$ reduces to the computation of the thermal one point functions of fermion bilinears from the $g_{3}(x)$ for the large $N$ Gross-Neveu model performed in David:2023uya , except the fact that one has to keep track of the $e^{im\mu}$ in the correlator (90) compared to the the correlator given in equation (4.33) of David:2023uya with d=3d=3(or $k=1$ as the notation used there d=2k+1d=2k+1) [^4] Note that there was a typo in the equation (4.33) of David:2023uya , in the denominator of this formula $(k+\frac{1}{2})_{l}$ is not correct, it should be replaced by $(k-\frac{1}{2})_{l}$.. And keeping this $e^{im\mu}$ in that calculation we get
@@
$).\displaystyle=\frac{1}{\pi(\frac{1}{2})_{l}}\Big{(}\frac{m_{f}}{2}\Big{)}^{l+2}\sum_{n=0}^{l}\frac{(l-n+1)_{2n}}{(2m_{f})^{n}n!}({\rm Li}_{n+1}(-e^{-m_{f}-i\hat{\mu}})+(-1)^{l}{\rm Li}_{n+1}(-e^{-m_{f}+i\hat{\mu}}$)). (91)
@@
For the one point function of the fermion bilinear ${\cal O}_{0}[0,0]$, the arc contribution has also been evaluated in David:2023uya and it is given by  [^6] See equation (4.49) in David:2023uya , note that the presence of the chemical potential $i\hat{\mu}$ doesn’t alter this answer, we kept track of an overall -ve sign in our correlator $g_{1}(z,\bar{z})$ in (79) compared to that used in David:2023uya .
@@ 6 Agreement with the stress tensor and spin-1 current
$\log Z$ $).\displaystyle=-N\beta V(\frac{m_{f}^{3}}{8\lambda}-\frac{m_{b}^{2}m_{f}}{8\lambda}-\frac{1}{\beta}\log Z_{b}(m_{b},-i\hat{\mu})-\frac{1}{\beta}\log Z_{f}(m_{f},-i\hat{\mu}$)). (117)
@@
Thus the combination $4a_{\mathcal{O}}[0,2]+2a_{\mathcal{O}_{+}}[0,2]$ agrees with the stress tensor for our model at the free CFT point where the coupling constant $\lambda=0$. We can also see that the expression (6) exactly matches with the stress tensor of the theory (6) at the interacting CFT obtained from the direct calculation of the partition function at $\beta=1$.
@@
$)\displaystyle=\frac{1}{{2\pi\beta^{2}}}\Big{[}\beta m_{b}(\log(1-e^{-\beta m_{b}+i\beta\hat{\mu}})-\log(1-e^{-\beta(m_{b}+i\hat{\mu})}$))
$)\displaystyle+\text{Li}_{2}(e^{-\beta(m_{b}+i\hat{\mu})})-\text{Li}_{2}(e^{i\beta\hat{\mu}-m_{b}\beta})+\beta m_{f}(\log(1+e^{-\beta(m_{f}+i\hat{\mu})})-\log(1+e^{-\beta m_{f}+i\beta\hat{\mu}}$))
@@ 7 Conclusions
We have studied the large $N$ Wess-Zumino model in presence of the chemical potential($\mu$) in $3d$. Starting from the supersymmetric Wess-Zumino model for $2N$ real scalars and $2N$ real fermions without chemical potential in the fundamental of $O(2N)$, we construct the $U(N)$ Wess-Zumino model with $N$ complex scalars and $N$ Dirac fermions. We have introduced a chemical potential in this setup at finite temperature. The non-trivial fixed point of the model is obtained at the saddle point of the action in the zero modes of the auxiliary fields used to linearise the action at large $N$. The saddle point condition is known as the gap equation relating the thermal masses to the chemical potential $\mu$ at finite temperature. These thermal masses, obtained by solving the gap equations, describe the non-trivial fixed point of the theory. The interaction between boson and fermion degrees of freedom in this model leads to a pair of coupled gap equations in this case.
@@
We have found the analytically solvable regime of the gap equations for the large $N$ Wess-Zumino model in $3d$ at large chemical potential($\mu$) for arbitrary coupling($\lambda$). By using these analytic solutions of the gap equation at large $\mu$, we evaluate the partition function, stress tensor and spin-1 current of the theory as a perturbative expansion at large $\mu$. Applying the OPE inversion formula we also evaluate the higher spin currents of the theory as a perturbative expansion at large $\mu$. As a direct consequence of the study of the OPE inversion formula we find that at the leading order in large $N$ the OPE coefficients $C_{\phi\phi{\cal O}_{0}[0,l]}$ and $C_{\psi\psi{\cal O}[0,l]}$ vanish for all values of $l$ and thus the relevant three point functions also vanish. Although our OPE analysis can not directly conclude whether the coefficients $C_{\phi\phi{\cal O}_{+}[0,l]}$ and $C_{\phi\phi{\cal O}_{-}[0,l]}$ also vanish. But comparing the results from the OPE inversion formula and the results obtained from the partition function we have shown that the $C_{\phi\phi{\cal O}_{+}[0,l]}$ indeed vanishes for $l=1,2$. It would be interesting to evaluate these coefficients directly.
An interesting problem would be to extend this study on curved backgrounds such as $S^{1}\times S^{2}$, the method developed in David:2024pir for $O(N)$ model in this geometry can be useful in this context. The large chemical potential limit may serve as a solvable limit in various other related models. The analysis presented in this paper perhaps contain useful ingredients necessary for the study of the large $N$ Gross-Neveu-Yukawa model Moshe:2003xn at finite chemical potential and finite temperature.
In the study of thermal one point functions of higher spin currents, large spin behaviour of these one point functions emerges as an interesting observationIliesiu:2018fao ; David:2023uya ; David:2024naf ; David:2024pir . For the large $N$ model of complex scalars and the Gross-Neveu model at finite chemical potential, the large spin limit of the 1-point functions of the currents at the non-trivial fixed point resembles the currents in free CFT. The dependence on the chemical potential($\mu$) reduces to a simple factor of $\sinh\mu$ or $\cosh\mu$ David:2024naf for odd or even spin respectively. Thus an interesting exploration would be to understand the interplay between large spin($l$) and large $\mu$ and to quantify it.
The study of the class of vector models at finite $N$ should be important for practical interests. For the $O(N)$ model at finite temperature and zero chemical potential, a bootstrap method Barrat:2024fwq is developed based on a set of sum rules Marchetto:2023xap . Using this method, thermal OPE coefficients are computed at finite $N$. It would be useful to generalise such a method in the presence of a finite chemical potential. The generalisation of the techniques of the OPE inversion formula for the large $N$ Chern-Simons matter theories Giombi:2011kc ; Aharony:2011jz still remains an interesting open question.